OpenStudy (anonymous):

find f'(g(x)) using the chain rule. y = (1+cot^5(x^4+1))^9

OpenStudy (accessdenied):

I'm confused... we'd use the chain rule multiple times, what "f'(g(x))" is it referring to...?

OpenStudy (anonymous):

its a chain within a chain within a chain...

OpenStudy (accessdenied):

y = (1+cot^5(x^4+1))^9 y = (1 + (cot(x^4 + 1))^5)^9 y' = 9(1 + (cot(x^4 + 1))^5)^8 * (0 + 5(cot(x^4 + 1))^4) * (-(csc(x^4 + 1))^2) * (4x^3) = -180x^3 (csc(x^4 + 1))^2 (cot(x^4 + 1))^4 (1 + (cot(x^4 + 1))^5)^8 A) f(x) = x^9 g(x) = 1 + (cot(x^4 + 1))^5 f'(x) = 9x^8 f'(g(x)) = 9(1 + (cot(x^4 + 1))^5)^8 B) f(x) = x^5 g(x) = cot(x^4 + 1) f'(x) = 5x^4 f'(g(x)) = 5(cot(x^4 + 1))^4 C) f(x) = cot(x) g(x) = x^4 + 1 f'(x) = -(csc(x))^2 f'(g(x)) = -(csc(x^4 + 1))^2